Solid lithium hydroxide is used in space vehicles to remove exhaled carbon dioxide from the living environment to form solid lithium carbonate and liquid water. What mass of gaseous carbon dioxide can be absorbed by 1.00kg of lithium hydroxide?

2LiOH(s) + CO2(g) → Li2CO3(s) + H2O(l)
1
Expert's answer
2020-05-10T14:57:34-0400

"2LiOH(s) + CO_2(g) \u2192 Li_2CO_3(s) + H_2O(l)"

we can see that ratio of stochiometric coefficients ratio of LiOH and carbon dioxide is 2:1

which means that for every two moles of LiOH , one mole of carbon dioxide is consumed


moles of LiOH "=\frac{1000}{24}=41.67"

moles of carbon dioxide consumed will be half of LiOH=41.67/2=20.83

mass of CO2 consumed =20.83"\times"44 =916.52 grams